OpenStudy (anonymous):

1.)solve the systems of equations using substitution y = 2x+3 y = 3x +1 2.)solve the system using elimination 3x + y =11 4x - y =-17 3.)5x + 8y =-29 7x - 2y=-67 4.)3x - y = 28 3x + y =14

Parth (parthkohli):

it can be said that 2x + 3 = 3x + 1 2x + 2 = 3x x = 2 Now that we have x, then let's find y. y = 2(2) + 3 y = 4 + 3 y = 7

Parth (parthkohli):

I follow "One question at a time" lol

OpenStudy (anonymous):

thank you lol i appreciate it now can you help me with the rest please haha :(